# Information-Driven Design of Imaging Systems

### v4

@@ Acknowledgements
The most effective approach to this challenge is to minimize system variations through careful engineering: building mechanically stable optical systems, using temperature control to reduce thermal drift, and implementing precise electronic control of components like illumination sources. However, some level of system variation often remains unavoidable in practice.
When measurements contain information about both the object and the system itself, the estimation of mutual information can be modified by considering the encoder state. Rather than estimating the full entropy of noisy measurements $H(\mathbf{Y})$, the conditional entropy $H(\mathbf{Y}|\mathbf{E})$ accounts for system variation, where $\mathbf{E}$ represents the encoder state. The conditional entropy estimate $H(\mathbf{Y}|\mathbf{X})$ remains unchanged, but $H(\mathbf{Y})$ includes randomness from three sources:
  • Object variations
  • System variations
  • Measurement noise
Conditioning on the encoder state $\mathbf{E}$ eliminates the contribution of system variations, isolating the object and noise entropy components.
In practice, conditioning can be accomplished by providing additional encoder state information like field position or calibration images to the entropy model during training. Flexible models like PixelCNN can readily accept such conditional data [46]. This reduces the entropy estimate by eliminating variations predictable from the system state.
We leave developing such conditional entropy estimation to future work. However, we can already observe the impact of encoder uncertainty in practical systems - in Section S4.1, we demonstrate how it affects information estimates in LED array microscopy.
@@
The likelihood of a probability distribution can be very sensitive to small changes in parameters because the Euclidean distance between parameter vectors does not always accurately reflect the dissimilarity of the resulting distributions (as described in section 2.3 of [113]). In theory this can be corrected by multiplying the gradient by the inverse of the Fisher information matrix to reorient it from the steepest direction in Euclidean space to the steepest direction in Riemannian space, which accounts for the natural geometry of the parameters. However, in practice we found this to be expensive to compute without obvious performance benefits over gradient clipping.
Figure S11 shows the results when fitting stationary Gaussian processes to images from the BSCCM dataset [65]. The direct estimate contained several negative eigenvalues, which were then set to an arbitrary eigenvalue floor. The optimization procedure was able to correct these negative eigenvalues, and produced robust results for a variety of settings of the eigenvalue floor parameter (provided it was smaller than the true minimum eigenvalue) (Fig. S11b). As a consequence of the small, incorrect eigenvalues in the direct estimate, samples produced from the model that were larger than the patch size on which they were trained (Sec. S2.6) exhibited numerical instability that created large oscillation in the samples. In contrast, samples from the optimized model were stable and did not exhibit this oscillation (Fig. S11b). The optimization quickly converged in a few iterations (Fig. S11c), taking $\sim$ 3 seconds to complete on an NVIDIA GeForce RTX 3090 GPU. Comparing estimates of a non-stationary Gaussian process, the direct estimate stationary Gaussian process, and the optimized stationary Gaussian process showed that the optimized stationary Gaussian process produced more robust results in terms of its eigenvalue distribution using smaller datasets (Fig. S11d).
@@
The comparable performance of stationary Gaussian and PixelCNN models on the BSCCM dataset (single cells under coded-angle illumination) likely reflects the near-Gaussian statistics of these images. However, this similarity cannot be assumed for all datasets. This becomes evident when testing on the MNIST handwritten digits dataset [114] (GNU General Public License), where image statistics deviate significantly from Gaussian distributions due to the nearly bimodal nature of pixel values (predominantly 0 or 1).
@@
Using simulated data derived from the BSCCM [65] dataset, we first examined the consistency of our conditional entropy estimator. We created ground truth data by applying a $3\times 3$ median filter to produce noiseless images, then added simulated shot noise. Our estimator rapidly converges to the true value (Fig. S18a), with variations on the order of $10^{-2}$ differential entropy per pixel - significantly smaller than the variations in $H(\mathbf{Y})$ estimation.
@@
We compared both approaches empirically across three imaging modalities in the BSCCM dataset using three estimators: analytic stationary Gaussian entropy, stationary Gaussian upper bound, and PixelCNN upper bound (Fig. S19). While we lack theoretical guarantees for the analytic stationary Gaussian estimates, they consistently track close to the PixelCNN bounds, which represent our best estimate of the true entropy. The differences between estimation methods remain small compared to the variations across imaging modalities and photon counts. This close agreement is particularly relevant for our IDEAL framework (See Information-Driven Encoder Analysis Learning (IDEAL)), where we use the analytic form in our optimization loss function.
@@
S4.1 Experimental evidence of encoder uncertainty effects
Figure S21: The effect of encoder uncertainty on information estimates. By computing mutual information estimates on empty image patches with no objects in them, the information about the system present in the measurements can be estimated. This can be seen by comparing Single LED, Differential Phase Contrast, and Brightfield illumination patterns on an LED array microscope, and estimating the mutual information between noiseless image and noisy measurement with and without an object. Encoders that capture more information about the object also capture more information about the system.
To demonstrate how encoder uncertainty (Sec. S1.5.1) affects information estimates in practice, we examined measurements from an LED array microscope using different illumination patterns. We estimated the mutual information in empty image patches containing no biological sample, allowing us to directly measure information content arising purely from system variations rather than objects of interest.
As shown in Figure S21, single-LED illumination exhibited higher background information content than either differential phase contrast or brightfield illumination. This aligns with the theoretical prediction that more sensitive imaging methods encode more information about both objects and system imperfections. The measurements reveal a fundamental tradeoff: illumination patterns that capture more object information also capture more system-derived information, potentially inflating information estimates.
These results highlight the practical importance of developing the conditional entropy estimation methods described in Section S2.7 to separate object information from system variations on highly sensitive systems with uncontrolled and unmodeled sources of randomness. Without such corrections, information estimates may be tend to overestimate the true information content for highly sensitive encoders or those with large amounts of uncontrollable system randomness.
@@
Mathematical assumptions about object structure underlie several established frameworks for imaging system design. Like our approach, estimation theory [115] and compressed sensing [116] offer ways to design imaging systems without requiring ground truth data or human-interpretable measurements. Estimation theory assumes objects can be described by a few key parameters and uses the Cramér-Rao lower bound [117, 86] to optimize system performance. Compressed sensing assumes objects are sparse in some basis, enabling efficient measurement with theoretical guarantees.
@@
Estimation theory provides tools for optimizing imaging systems by minimizing measurement uncertainty. Its central tool, the Cramér-Rao lower bound [115], defines the minimum achievable error when estimating object parameters from noisy measurements. This bound enables systematic comparison of different optical designs by identifying those with the lowest theoretical error.
@@
The Cramér-Rao lower bound has guided the design of numerous imaging systems. Its most successful application has been in fluorescence microscopy, where it has helped optimize point spread functions for localizing single molecules [118, 117, 119, 120, 121, 86, 122]. Similar approaches have proven valuable in phase microscopy [123] and wavefront sensing [124].
@@
It’s important to note that Fisher information, despite sharing the word “information,” is distinct from Shannon’s mutual information and entropy used elsewhere in this work. While both concepts quantify aspects of measurement uncertainty, they arose from different fields—estimation theory and information theory—and were developed independently. Though theoretical connections exist between them in certain cases [115], they serve different purposes in analysis and design.
@@
The approach of designing physical systems based on the Cramér-Rao lower bound has inherent limitations that prevent its application outside of a narrow set of applications in which the decoding problem is limited to estimation of one or a few parameters. Furthermore, even when it can be applied, it requires developing a complete mathematical model of the physics of image formation with respect to the parameter of interest. These limitations have thus far prevented its application outside a narrow range of problems, which include estimation of the position or mass of a single particle [118, 117, 119, 120, 121, 86, 125], or estimating a single optical aberration such as defocus [124]. Below, we discuss these limitations in more detail.
@@
However, creating a straightforward equation-based model for imaging most types of objects is usually very challenging due to the difficulty and increasing complexity of adding additional parameters. Most objects being imaged lack the simplicity of a single point emitter, which can be fully described by its XYZ location. One potential way to circumvent this issue is a more general purpose parameterization of the object in which it is represented by a discrete array of pixel values [123]. However, unbiased estimators are rarely used on high-dimensional estimation problems like this. Furthermore, the Cramér-Rao lower bound is in terms of mean squared error, which, when applied to an array of pixels, generally does not effectively capture semantically meaningful information about objects [126].
@@
In light of the the difficulties of extending this approach to objects with high-dimensional parameterizations, another possibility is to design imaging systems on simple classes of objects with the hope that they will generalize to other classes of objects. Taking such an approach requires making additional assumptions with unknown effects on results. Our experiments in minimal 1-Dimensional simulations show that the object-dependence of imaging systems can be readily demonstrated (Sec. S1.3), which may in part explain why empirical solutions to particle localization problems deviate from theoretical predictions [122].
@@
In addition to the practical difficulties of formulating complex parametric models and computing corresponding Cramér-Rao lower bounds, there remains the limitation that the (standard) Cramér-Rao lower bound quantifies the minimum mean squared error only of unbiased estimators. While constraining an estimator to be unbiased is arguably a reasonable choice for simple, low-dimensional parameters like the location of a single point in 3D space, most estimators used in practice are in fact biased [127, 128] and state-of-the-art methods for solving image processing tasks almost exclusively use biased estimators to achieve high performance.
The bias-variance tradeoff [129] provides a useful perspective: biased estimators, while possessing higher bias, can have lower variance, thereby potentially reducing overall error. Thus, with bias, estimators with better error than the standard Cramér-Rao lower bound can be achieved. A classic illustration of this principle is the James-Stein estimator [130], which improves the estimation of multiple parameters simultaneously by shrinking individual estimates towards a common mean. It is based on the counterintuitive principle that, under certain conditions, an estimator that partially pools the data towards a central value can produce overall estimates that are closer to the true values than those obtained by estimating each parameter independently, especially when dealing with small sample sizes or high-dimensional data. It results in lower average error than the unbiased sample mean approach for Gaussian random variables when the number of dimensions is $\geq 3$.
Furthermore, the advantages of biased estimators are clear on empirical problems, particularly high-dimensional ones. State of the art methods on image-to-image estimation problems like denoising and deconvolution are usually achieved using deep neural networks [131] or with iterative optimization procedures that use regularization to bias the estimates towards certain classes of solutions.
@@
Like the standard Cramér-Rao lower bound, the van Trees inequality (also known as the Bayesian Cramér-Rao lower bound) [132, 127] provides a lower bound on the squared error that can be achieved in a parameter estimation problem. Mathematically:
@@
Originally, information-theoretic quantities like entropy and mutual information were developed for noise-affected message transmission. However, these tools have significant theoretical links to estimations of the precision of noise-corrupted random variables. Though these connections are mostly known only for simpler analytical cases such as Gaussian random variables, they nonetheless provide insights into the relationships between these fields [115]. Here we highlight some of these connections.
@@
Much of estimation theory centers on limits defined in terms of mean squared error of signals. While mean squared error has many appealing properties, its shortcomings, particularly in the context of quantifying the perceptual and semantic quality of images are readily apparent [126]. This has, for example, motivated work on alternative ways of quantifying error [133, 134].
@@
In contrast, rate distortion theory, a branch of information theory, can be used to understand the fundamental limits and behavior of a wide variety of loss functions. The connection from the lower bounds used in estimation theory and information theory can be readily seen in the Efroimovich inequality [135], which generalizes the van Trees inequality (Equation 49) from providing a bound on only mean squared error to providing a bound on a more general way of quantifying uncertainty, the entropy of a parameter $\theta$ given a noisy measurement $X$:
@@
Here, $h(\Theta\mid X)$ is the differential entropy of a parameter given data $X$. This inequality can be used to derive bounds on loss functions other than mean squared error in terms of information-theoretic quantities [128, 136].
@@
Another known relationship with particular relevance to this work is known as the I-MMSE formula (short for Information - minimum mean squared error), which states that [137, 115]:
[…]
@@
This formula shows that mutual information quantifies the same operational idea as the (Bayesian) Cramér-Rao lower bound in the case of additive Gaussian noise, which strongly suggests that the quantities may serve similar purposes under more general noise models. This relationship has been leveraged to develop task-specific information estimators [23].
@@
When these assumptions hold, compressed sensing approaches theoretical performance limits [138]. However, real-world imaging often violates these assumptions. Natural objects rarely exhibit perfect sparsity, and measurement noise may not be Gaussian. Studies have shown that compressed sensing strategies become suboptimal when objects don’t match the assumed structure [31, 30].
Recent work has shown that incorporating additional object knowledge can improve upon basic compressed sensing. When signals can be generated by passing low-dimensional vectors through nonlinear models, reconstruction accuracy can improve by an order of magnitude compared to traditional compressed sensing approaches [139, 140, 141].
More broadly, probabilistic models offer a framework for describing signal structure beyond sparsity. These models enable analysis of average-case performance and can incorporate various forms of prior knowledge about signals [142, 143]. Theoretical work has shown how measurement limits depend fundamentally on the complexity of the objects being measured [144], and how probabilistic approaches can achieve optimal performance for specific signal classes [145, 146].
@@
The images were modeled using a PixelCNN architecture adapted from prior work [45, 46]. The model consists of vertical and horizontal masked convolution layers arranged in a stack, enabling modeling of the conditional distribution of each pixel given the previous pixels in raster scan order. The network includes seven gated masked convolution layers with alternating dilation rates (1,2,1,4,1,2,1) to increase the receptive field while maintaining computational efficiency. Each gated layer combines information from vertical and horizontal stacks through masked convolutions followed by a gating mechanism using tanh and sigmoid activations.
We modified the original categorical output distribution to instead use a Gaussian mixture density [111] with 40 components at each pixel. The mixture parameters (means, standard deviations, and mixing weights) were computed through separate dense layers. The Gaussian means were initialized randomly uniformly between the training image minimum and maximum values and clipped to this range during inference. The standard deviations were computed as the softplus output of a dense layer, clipped to be between 1 and the training set standard deviation to avoid degenerate solutions.

### v5

@@ Acknowledgements
The most effective approach to reduce encoder uncertainty is to minimize system variations through careful engineering: building mechanically stable optical systems, using temperature control to reduce thermal drift, and implementing precise electronic control of components like illumination sources. Nonetheless, some level of system variation often remains unavoidable in practice.
Here we describe how our framework can be expanded to stochastic encoders to account for encoder uncertainty.
S2.7.1 Encoder uncertainty
Stochastic encoders affect information estimates two ways. First, different encoder states preserve different amounts of object information. For example, defocus destroys sharp features in images. Average object information thus depends on the focus distribution.
Second, encoder variations create measurement diversity beyond object variations. With focus drift, measurements vary in both content and sharpness. Without modification, our method would interpret both as object information, thereby upwardly biasing its estimates.
One potential way to adapt our method is using calibration data to identify encoder-induced variations. This would likely require either additional compute or simplifying assumptions to handle the fact the the measurement diversity induced by encoder variations may be object dependent. Mathematical formulation
We model encoder variations with state variable $\mathbf{E}$:
$(\mathbf{O},\mathbf{E})\to\mathbf{X}\to\mathbf{Y}$
Now $\mathbf{X}=f(\mathbf{O},\mathbf{E})$ for some deterministic function $f$. For example, for focus drift, $f$ would describe how defocus $\mathbf{E}$ blurs object $\mathbf{O}$ to produce image $\mathbf{X}$.
The quantity of interest in the original deterministic setting is now written with the encoder explicit $I(\mathbf{O};\mathbf{Y}|\mathbf{e})$.
In the stochastic case, averaging over all encoder states gives:
$\bar{I}(\mathbf{O};\mathbf{Y})=\sum_{\mathbf{e}}p(\mathbf{e})I(\mathbf{O};\mathbf{Y}|\mathbf{e})$
Where $\bar{I}(\mathbf{O};\mathbf{Y})$ denotes the average object information across all encoder states.
Extending the deterministic information transfer theorem to $\mathbf{X}=f(\mathbf{O},\mathbf{E})$:
$I(\mathbf{X};\mathbf{Y})=I(\mathbf{O},\mathbf{E};\mathbf{Y})$
Rewriting using the chain rule of mutual information:
$I(\mathbf{X};\mathbf{Y})=\bar{I}(\mathbf{O};\mathbf{Y})+I(\mathbf{E};\mathbf{Y}|\mathbf{O})$
The first term captures effect one (average object information across encoder states); the second captures effect two (diversity from encoder variations).
Example: random focal drift
Consider a microscope with mechanical drift causing focus to vary between perfect focus and various degrees of defocus. Compared to a stable, in-focus system, we have $\bar{I}(\mathbf{O};\mathbf{Y})<I(\mathbf{O};\mathbf{Y}|\mathbf{e}_{\text{in-focus}})$ because defocus destroys high-frequency information. Meanwhile, $I(\mathbf{E};\mathbf{Y}\mid\mathbf{O})>0$ because focus variations add to measurement diversity.
As shown in the Deterministic Information Transfer Theorem, a deterministic in-focus system has $I(\mathbf{X};\mathbf{Y})=I(\mathbf{O};\mathbf{Y}|\mathbf{e}_{\text{in-focus}})$. With stochastic focus variations, $I(\mathbf{X};\mathbf{Y})$ can increase or decrease relative to this baseline depending on which of the two effects dominates.
S2.7.2 Estimating information with stochastic encoders
When encoders vary stochastically, our method estimates $I(\mathbf{X};\mathbf{Y})=\bar{I}(\mathbf{O};\mathbf{Y})+I(\mathbf{E};\mathbf{Y}|\mathbf{O})$. To isolate object information $\bar{I}(\mathbf{O};\mathbf{Y})$, we must estimate and subtract the encoder contribution.
The central challenge is that $I(\mathbf{E};\mathbf{Y}|\mathbf{O})$ depends on the object—different objects may reveal encoder variations differently. We propose two approaches:
Approach 1: Fix object, vary encoder states, estimate $I(\mathbf{E};\mathbf{Y}|\mathbf{O}=\mathbf{o})$ for multiple objects. This would be feasible in IDEAL using increased compute, but in experimental systems only possible if we can fix objects while varying encoder states (e.g., imaging a static sample while inducing focus drift).
Section S4.1 shows an example of non-object information in the measurement.
Approach 2: Assume $I(\mathbf{E};\mathbf{Y}|\mathbf{O})\approx I(\mathbf{E};\mathbf{Y})$. Computationally and experimentally simpler, but may give incorrect results if encoder effects are highly object-dependent.
Taking Approach 2 as an example:
$I(\mathbf{E};\mathbf{Y})=H(\mathbf{Y})-H(\mathbf{Y}|\mathbf{E})$
Our method already estimates $H(\mathbf{Y})$. To estimate $H(\mathbf{Y}|\mathbf{E})$, we fit conditional models $p_{\theta}(\mathbf{y}|\mathbf{e})$ instead of $p_{\theta}(\mathbf{y})$, where $\mathbf{e}$ represents encoder state information. For focus drift, conditioning on timestamps would help predict measurements since focus varies smoothly with time.
For Approach 1, we would additionally need to take an outer expectation over objects. In IDEAL, since we know the true object, the models could be modified to condition on it directly. In experimental conditions, we would likely need static objects to take multiple measurements of.
By estimating both terms we can recover the quantity of interest, object information in the measurement.
$\bar{I}(\mathbf{O};\mathbf{Y})=I(\mathbf{X};\mathbf{Y})-I(\mathbf{E};\mathbf{Y}|\mathbf{O})$
S2.8 Task-specific information
While information estimates strongly predict reconstruction performance, they may correlate less strongly with specialized tasks that rely only on specific features of measurements. Systems that preserve specific task-relevant information can outperform those capturing more total information indiscriminately [neifeldTaskspecificInformationImaging2007, ashokCompressiveImagingSystem2008]
Total information naturally decomposes into task-specific and task-irrelevant components: $I(\mathbf{Y};\mathbf{O})=I(\mathbf{Y};\mathbf{T})+I(\mathbf{Y};\mathbf{O}|\mathbf{T})$, where $\mathbf{T}$ represents task variables like classification labels. This enables weighted objectives $\alpha I(\mathbf{Y};\mathbf{T})+\beta I(\mathbf{Y};\mathbf{O}|\mathbf{T})$ for specialized applications:
  • Balanced systems ($\alpha,\beta>0$): Control relative importance of task-specific versus general information
  • Compressed sensing ($\alpha>0,\beta<0$): Maximize task-relevant while suppressing task-irrelevant information for bandwidth-limited applications
  • Privacy-preserving ($\alpha<0,\beta>0$): Suppress task-specific information while preserving general scene information
Estimation requires computing $I(\mathbf{Y};\mathbf{T})=H(\mathbf{Y})-H(\mathbf{Y}|\mathbf{T})$. Our existing approach estimates $H(\mathbf{Y})$, while $H(\mathbf{Y}|\mathbf{T})$ can be estimated by conditioning models on task information. Section S4.2 provides experimental evidence of when information estimates do not fully predict task performance.
@@
The likelihood of a probability distribution can be very sensitive to small changes in parameters because the Euclidean distance between parameter vectors does not always accurately reflect the dissimilarity of the resulting distributions (as described in section 2.3 of [hoffmanStochasticVariationalInference2013]). In theory this can be corrected by multiplying the gradient by the inverse of the Fisher information matrix to reorient it from the steepest direction in Euclidean space to the steepest direction in Riemannian space, which accounts for the natural geometry of the parameters. However, in practice we found this to be expensive to compute without obvious performance benefits over gradient clipping.
Figure S11 shows the results when fitting stationary Gaussian processes to images from the BSCCM dataset [pinkardBerkeleySingleCell2024]. The direct estimate contained several negative eigenvalues, which were then set to an arbitrary eigenvalue floor. The optimization procedure was able to correct these negative eigenvalues, and produced robust results for a variety of settings of the eigenvalue floor parameter (provided it was smaller than the true minimum eigenvalue) (Fig. S11b). As a consequence of the small, incorrect eigenvalues in the direct estimate, samples produced from the model that were larger than the patch size on which they were trained (Sec. S2.6) exhibited numerical instability that created large oscillation in the samples. In contrast, samples from the optimized model were stable and did not exhibit this oscillation (Fig. S11b). The optimization quickly converged in a few iterations (Fig. S11c), taking $\sim$ 3 seconds to complete on an NVIDIA GeForce RTX 3090 GPU. Comparing estimates of a non-stationary Gaussian process, the direct estimate stationary Gaussian process, and the optimized stationary Gaussian process showed that the optimized stationary Gaussian process produced more robust results in terms of its eigenvalue distribution using smaller datasets (Fig. S11d).
@@
The comparable performance of stationary Gaussian and PixelCNN models on the BSCCM dataset (single cells under coded-angle illumination) likely reflects the near-Gaussian statistics of these images. However, this similarity cannot be assumed for all datasets. This becomes evident when testing on the MNIST handwritten digits dataset [lecunGradientbasedLearningApplied1998] (GNU General Public License), where image statistics deviate significantly from Gaussian distributions due to the nearly bimodal nature of pixel values (predominantly 0 or 1).
@@
Using simulated data derived from the BSCCM [pinkardBerkeleySingleCell2024] dataset, we first examined the consistency of our conditional entropy estimator. We created ground truth data by applying a $3\times 3$ median filter to produce noiseless images, then added simulated shot noise. Our estimator rapidly converges to the true value (Fig. S18a), with variations on the order of $10^{-2}$ differential entropy per pixel - significantly smaller than the variations in $H(\mathbf{Y})$ estimation.
@@
We compared both approaches empirically across three imaging modalities in the BSCCM dataset using three estimators: analytic stationary Gaussian entropy, stationary Gaussian upper bound, and PixelCNN upper bound (Fig. S19). While we lack theoretical guarantees for the analytic stationary Gaussian estimates, they consistently track close to the PixelCNN bounds, which represent our best estimate of the true entropy. The differences between estimation methods remain small compared to the variations across imaging modalities and photon counts. This close agreement is particularly relevant for our IDEAL framework (See Encoder design via information maximization with IDEAL), where we use the analytic form in our optimization loss function.
@@
S4.1 Experimental evidence of stochastic encoder effects
Figure S21: The effect of stochastic encoders on information estimates. By computing mutual information estimates on empty image patches with no objects in them, the information about the system present in the measurements can be estimated. This can be seen by comparing Single LED, Differential Phase Contrast, and Brightfield illumination patterns on an LED array microscope, and estimating the mutual information between noiseless image and noisy measurement with and without an object. Encoders that capture more information about the object also capture more information about the system.
We examined LED array microscopy with different illumination patterns to demonstrate how sensitive systems capture both object and system information. By estimating mutual information on empty image patches, we measured information arising purely from system variations.
As shown in Figure S21, single-LED illumination exhibited higher background information than differential phase contrast or brightfield illumination. This demonstrates that more sensitive encoders capture more system information alongside object information. When substantial encoder uncertainty is present, the theoretical framework from Section S2.7 may be needed to separate these contributions.
@@
Mathematical assumptions about object structure underlie several established frameworks for imaging system design. Like our approach, estimation theory [guoInterplayInformationEstimation2013] and compressed sensing [candesIntroductionCompressiveSampling2008] offer ways to design imaging systems without requiring ground truth data or human-interpretable measurements. Estimation theory assumes objects can be described by a few key parameters and uses the Cramér-Rao lower bound [pavaniHighefficiencyRotatingPoint2008, chaoFisherInformationTheory2016] to optimize system performance. Compressed sensing assumes objects are sparse in some basis, enabling efficient measurement with theoretical guarantees.
@@
Estimation theory provides tools for optimizing imaging systems by minimizing measurement uncertainty. Its central tool, the Cramér-Rao lower bound [guoInterplayInformationEstimation2013], defines the minimum achievable error when estimating object parameters from noisy measurements. This bound enables systematic comparison of different optical designs by identifying those with the lowest theoretical error.
@@
The Cramér-Rao lower bound has guided the design of numerous imaging systems. Its most successful application has been in fluorescence microscopy, where it has helped optimize point spread functions for localizing single molecules [pavaniThreeDimensionalTracking2008a, pavaniHighefficiencyRotatingPoint2008, shechtmanOptimalPointSpread2014, groverPerformanceLimitsThreedimensional2010, quirinOptimal3DSinglemolecule2012, chaoFisherInformationTheory2016, nehmeDeepSTORM3DDense3D2020]. Similar approaches have proven valuable in phase microscopy [bouchetFundamentalBoundsPrecision2021] and wavefront sensing [leeCramerRaoAnalysis1999a].
@@
It’s important to note that Fisher information, despite sharing the word “information,” is distinct from Shannon’s mutual information and entropy used elsewhere in this work. While both concepts quantify aspects of measurement uncertainty, they arose from different fields—estimation theory and information theory—and were developed independently. Though theoretical connections exist between them in certain cases [guoInterplayInformationEstimation2013], they serve different purposes in analysis and design.
@@
The approach of designing physical systems based on the Cramér-Rao lower bound has inherent limitations that prevent its application outside of a narrow set of applications in which the decoding problem is limited to estimation of one or a few parameters. Furthermore, even when it can be applied, it requires developing a complete mathematical model of the physics of image formation with respect to the parameter of interest. These limitations have thus far prevented its application outside a narrow range of problems, which include estimation of the position or mass of a single particle [pavaniThreeDimensionalTracking2008a, pavaniHighefficiencyRotatingPoint2008, shechtmanOptimalPointSpread2014, groverPerformanceLimitsThreedimensional2010, quirinOptimal3DSinglemolecule2012, chaoFisherInformationTheory2016, dongFundamentalBoundsPrecision2021], or estimating a single optical aberration such as defocus [leeCramerRaoAnalysis1999a]. Below, we discuss these limitations in more detail.
@@
However, creating a straightforward equation-based model for imaging most types of objects is usually very challenging due to the difficulty and increasing complexity of adding additional parameters. Most objects being imaged lack the simplicity of a single point emitter, which can be fully described by its XYZ location. One potential way to circumvent this issue is a more general purpose parameterization of the object in which it is represented by a discrete array of pixel values [bouchetFundamentalBoundsPrecision2021]. However, unbiased estimators are rarely used on high-dimensional estimation problems like this. Furthermore, the Cramér-Rao lower bound is in terms of mean squared error, which, when applied to an array of pixels, generally does not effectively capture semantically meaningful information about objects [zhouwangMeanSquaredError2009].
@@
In light of the the difficulties of extending this approach to objects with high-dimensional parameterizations, another possibility is to design imaging systems on simple classes of objects with the hope that they will generalize to other classes of objects. Taking such an approach requires making additional assumptions with unknown effects on results. Our experiments in minimal 1-Dimensional simulations show that the object-dependence of imaging systems can be readily demonstrated  (Sec. S1.3), which may in part explain why empirical solutions to particle localization problems deviate from theoretical predictions [nehmeDeepSTORM3DDense3D2020].
@@
In addition to the practical difficulties of formulating complex parametric models and computing corresponding Cramér-Rao lower bounds, there remains the limitation that the (standard) Cramér-Rao lower bound quantifies the minimum mean squared error only of unbiased estimators. While constraining an estimator to be unbiased is arguably a reasonable choice for simple, low-dimensional parameters like the location of a single point in 3D space, most estimators used in practice are in fact biased [gillApplicationsVanTrees1995, arasFamilyBayesianCram2019] and state-of-the-art methods for solving image processing tasks almost exclusively use biased estimators to achieve high performance.
The bias-variance tradeoff [hastieElementsStatisticalLearning2009] provides a useful perspective: biased estimators, while possessing higher bias, can have lower variance, thereby potentially reducing overall error. Thus, with bias, estimators with better error than the standard Cramér-Rao lower bound can be achieved. A classic illustration of this principle is the James-Stein estimator [jamesEstimationQuadraticLoss1992], which improves the estimation of multiple parameters simultaneously by shrinking individual estimates towards a common mean. It is based on the counterintuitive principle that, under certain conditions, an estimator that partially pools the data towards a central value can produce overall estimates that are closer to the true values than those obtained by estimating each parameter independently, especially when dealing with small sample sizes or high-dimensional data. It results in lower average error than the unbiased sample mean approach for Gaussian random variables when the number of dimensions is $\geq 3$.
Furthermore, the advantages of biased estimators are clear on empirical problems, particularly high-dimensional ones. State of the art methods on image-to-image estimation problems like denoising and deconvolution are usually achieved using deep neural networks [lecunDeepLearning2015] or with iterative optimization procedures that use regularization to bias the estimates towards certain classes of solutions.
@@
Like the standard Cramér-Rao lower bound, the van Trees inequality (also known as the Bayesian Cramér-Rao lower bound) [vantreesDetectionEstimationModulation2001, gillApplicationsVanTrees1995] provides a lower bound on the squared error that can be achieved in a parameter estimation problem. Mathematically:
@@
Originally, information-theoretic quantities like entropy and mutual information were developed for noise-affected message transmission. However, these tools have significant theoretical links to estimations of the precision of noise-corrupted random variables. Though these connections are mostly known only for simpler analytical cases such as Gaussian random variables, they nonetheless provide insights into the relationships between these fields [guoInterplayInformationEstimation2013]. Here we highlight some of these connections.
@@
Much of estimation theory centers on limits defined in terms of mean squared error of signals. While mean squared error has many appealing properties, its shortcomings, particularly in the context of quantifying the perceptual and semantic quality of images are readily apparent [zhouwangMeanSquaredError2009]. This has, for example, motivated work on alternative ways of quantifying error [wangImageQualityAssessment2004, zhangUnreasonableEffectivenessDeep2018].
@@
In contrast, rate distortion theory, a branch of information theory, can be used to understand the fundamental limits and behavior of a wide variety of loss functions. The connection from the lower bounds used in estimation theory and information theory can be readily seen in the Efroimovich inequality [efroimovich1980information], which generalizes the van Trees inequality (Equation 50) from providing a bound on only mean squared error to providing a bound on a more general way of quantifying uncertainty, the entropy of a parameter $\theta$ given a noisy measurement $X$:
@@
Here, $h(\Theta\mid X)$ is the differential entropy of a parameter given data $X$. This inequality can be used to derive bounds on loss functions other than mean squared error in terms of information-theoretic quantities [arasFamilyBayesianCram2019, lee2022new].
@@
Another known relationship with particular relevance to this work is known as the I-MMSE formula (short for Information - minimum mean squared error), which states that [guoMutualInformationMinimum2005, guoInterplayInformationEstimation2013]:
$I(X;Y)=I(X;\sqrt{s}X+N)=\frac{1}{2}\int_{0}^{s}\operatorname{mmse}(X\mid\gamma X+N)d\gamma$
@@
This formula shows that mutual information quantifies the same operational idea as the (Bayesian) Cramér-Rao lower bound in the case of additive Gaussian noise, which strongly suggests that the quantities may serve similar purposes under more general noise models. This relationship has been leveraged to develop task-specific information estimators [neifeldTaskspecificInformationImaging2007].
@@
When these assumptions hold, compressed sensing approaches theoretical performance limits [candesHowWellCan2013]. However, real-world imaging often violates these assumptions. Natural objects rarely exhibit perfect sparsity, and measurement noise may not be Gaussian. Studies have shown that compressed sensing strategies become suboptimal when objects don’t match the assumed structure [changInformativeSensing2009, ashokInformationOptimalCompressive2014].
Recent work has shown that incorporating additional object knowledge can improve upon basic compressed sensing. When signals can be generated by passing low-dimensional vectors through nonlinear models, reconstruction accuracy can improve by an order of magnitude compared to traditional compressed sensing approaches [boraCompressedSensingUsing2017, weiStatisticalRateNonlinear, mixonSUNLayerStableDenoising2018].
More broadly, probabilistic models offer a framework for describing signal structure beyond sparsity. These models enable analysis of average-case performance and can incorporate various forms of prior knowledge about signals [baraniukModelBasedCompressiveSensing2009, wainwrightInformationtheoreticLimitsSparsity2007]. Theoretical work has shown how measurement limits depend fundamentally on the complexity of the objects being measured [wuRenyiInformationDimension2010], and how probabilistic approaches can achieve optimal performance for specific signal classes [donohoInformationTheoreticallyOptimalCompressed2013, reevesSamplingRateDistortionTradeoff2012].
@@
The images were modeled using a PixelCNN architecture adapted from prior work [oordPixelRecurrentNeural2016, vandenoordConditionalImageGeneration2016]. The model consists of vertical and horizontal masked convolution layers arranged in a stack, enabling modeling of the conditional distribution of each pixel given the previous pixels in raster scan order. The network includes seven gated masked convolution layers with alternating dilation rates (1,2,1,4,1,2,1) to increase the receptive field while maintaining computational efficiency. Each gated layer combines information from vertical and horizontal stacks through masked convolutions followed by a gating mechanism using tanh and sigmoid activations.
We modified the original categorical output distribution to instead use a Gaussian mixture density [bishopMixtureDensityNetworks1994] with 40 components at each pixel. The mixture parameters (means, standard deviations, and mixing weights) were computed through separate dense layers. The Gaussian means were initialized randomly uniformly between the training image minimum and maximum values and clipped to this range during inference. The standard deviations were computed as the softplus output of a dense layer, clipped to be between 1 and the training set standard deviation to avoid degenerate solutions.